A lot consists of 10 good articles, 4 with minor defects and 2 with major defects.




i. One article is chosen at random. Find the probability that




a. It has no defects, c) It is either good or has major defects.




b. It has no major defects,

1
Expert's answer
2022-03-22T04:13:58-0400

a. "P(no\; defects)=\frac{10-4-2}{10}=\frac{2}{5}=0.4."


b. "P(no\;major\;defects)=\frac{10-2}{10}=\frac{4}{5}=0.8."


c. "P(good\;or\;major \;defects)=\frac{10-4}{10}=\frac{3}{5}=0.6."